PQ = 6x + 25 and QR = 16 – 3x; Find PR. 


1
Expert's answer
2020-09-14T15:55:39-0400

"PR=\sqrt{(6x+25)²+(16-3x)²-2(6x+25)(16-3x)cos(\angle PQR)}=\sqrt{36x²+300x+625+256-96x+9x²-2(96x-18x²+400-75x)cos(\angle PQR)}=\sqrt{45x²+204x+881-2(21x-18x²+400)cos(\angle PQR)}.
Answer: PR=\sqrt{45x²+204x+881+2cos(\angle PQR)(18x²-21x-400)}"